OpenStudy (anonymous):

The slope of a line is -9 and the point (-10, -8) is on the line. Express the line in slope-intercept form.

OpenStudy (amistre64):

y = -9(x--10)-8 y = -9x -90 -8 y = -9-98

OpenStudy (amistre64):

i lost my x lol

OpenStudy (amistre64):

if anyone finds it; please return it to lost and found....

OpenStudy (anonymous):

lol, you are really smart, i could use ur help outside of this